The perimeter of a teacher's table is 4.8 meters. The length is 60 cm shorter than twice the width. Find the dimensions of the table.
Answer by ikleyn(52778)

.

Let the width be x meters.

Then the length is (2x-0.6) meters.


Write an equation for the perimeter

    x + (2x-0.6) + x + (2x-0.6) = 4.8  meters.


Simplify and find x

    6x - 1.2 = 4.8

    6x = 4.8 + 1.2

    6x =     6.0

     x = 1.


ANSWER.  The width is 1 meter.  The length is 2x-0.6 = 2*1-0.6 = 1.4  meters.


CHECK.  The perimeter is  1 + 1.4 + 1 + 1.4 = 4.8 meters.  ! correct !

Solved.